# Identification of Single-Treatment Effects in Factorial Experiments

### v2

@@ Abstract
Despite their cost, randomized controlled trials (RCTs) are widely regarded as gold-standard evidence in disciplines ranging from social science to medicine. In recent decades, researchers have increasingly sought to reduce the resource burden of repeated RCTs with factorial designs that simultaneously test multiple hypotheses, e.g. experiments that evaluate the effects of many medications or products simultaneously. Here I show that when multiple interventions are randomized in experiments, the effect any single intervention would have outside the experimental setting is not identified absent heroic assumptions, even if otherwise perfectly realistic conditions are achieved. This happens because single-treatment effects involve a counterfactual world with a single focal intervention, allowing other variables to take their natural values (which may be confounded or modified by the focal intervention). In contrast, observational studies and factorial experiments provide information about potential-outcome distributions with zero and multiple interventions, respectively. In this paper, I formalize sufficient conditions for the identifiability of those isolated quantities. I show that researchers who rely on this type of design have to justify either linearity of functional forms or— in the nonparametric case — specify with Directed Acyclic Graphs how variables are related in the real world. Finally, I develop nonparametric sharp bounds—i.e., maximally informative best-/worst-case estimates consistent with limited RCT data—that show when extrapolations about effect signs are empirically justified. These new results are illustrated with simulated data.
Keywords: Factorial Experiment; Partial identification; Causal inference; Confounding
@@ 1 Introduction
Randomized controlled trials are now widely considered the gold standard in empirical research. Researchers have utilized them to answer questions ranging from the effects of medication to economic development and discrimination. But their logistical complexity and cost are compounded by scientists’ common need to test many hypotheses. As a result, scholars turn to factorial experiments, which evaluate multiple hypotheses simultaneously. Papers using this general design number in the tens of thousands [^1] A Google Scholar search in December 2023 returned 94,500 results for “factorial experiment”. Recent theoretical results were reported by Dasgupta et al., (2015); Lee et al., (2020); Lee and Shpitser, (2020); Egami and Imai, (2018); De la Cuesta et al., (2022); Zhao and Ding, (2022); Blackwell and Pashley, (2023); Pashley and Bind, (2023). and variants like conjoint analysis are an increasingly dominant approach in areas like market research and survey experiments. Among the advantages of that design, statisticians believe they provide greater efficiency because they allow the evaluation of factors with only one-quarter of the data that otherwise would be necessary, and greater comprehensiveness, as not only a few factors are analyzed but also their interactions (Fisher,, 1949, p. 113). Recently, Nobel of Economics laureates Esther Duflo and Michael Kremer, argued, “because they significantly reduce costs, cross-cutting different treatments [i.e., factorial design] has proved very important in allowing for the recent wave of randomized evaluations in development economics” (Duflo et al.,, 2007). It is widely believed that randomization allows scientists to infer “model-free”, or nonparametric, estimates of average treatment effects.
@@
Contrary to this common belief, I show that in general, factorial experiments do not produce nonparametric causal inference of certain single-treatment quantities and require additional functional assumptions. Intuitively, this is because factorial experiments present a setting where multiple treatments are implemented at the same time, but when researchers are interested in single-treatment quantities, they are devising environments where one treatment is manipulated but all the others might be confounded with the outcome, i.e. other variables might be causing the treatment variable and the outcome at the same time. As those cases can only be assessed in trials with single treatments, any extrapolation requires identifiability proofs, which – I will show – do not exist for the total nonparametric case.
@@
Consider the following example. Suppose one is investigating individuals admitted to a hospital with symptoms of COVID-19. Let a binary $Y$ denote survival after 2 months. Assume $A$ denotes hydroxychloroquine, a drug often employed to treat lupus, and $B$ denotes admission or not to the Intensive Care Unit. Both are binary variables. One would like to know not only the joint effect of the two treatments, $A$ and $B$, on the outcome $Y$ but also their effects in isolation, for instance, the single-treatment effect of $A$ on $Y$ or $B$ on $Y$. If one designs a study, there are 9 cases we could consider: 1) 4 groups with 2-treatment interventions, A=B=1, A=0/B=1, A=1/B=0, and A=B=0; 2) 4 groups with 1-treatment interventions, A=0, A=1, B=0, B=1; 3) 1 group with no intervention, i.e. observational data. All those cases are represented in Figure 1. The first case is exactly the case of a factorial experiment, while the second case is a standard randomized controlled trial (RCT), and the third represents observational data. Notice that observational data is not the same thing as the case $A=0;B=0$ in the factorial experiment, because $A$ and $B$ might be confounded with $Y$ in the no-treatment group. Indeed those cases are not even similar to groups $B=0$ and $A=0$ in the Standard RCT, because if one intervenes on $A$, $B$ might still be confounded, and vice-versa, if one intervenes on $B$. Therefore, if one intends to draw inferences on the average treatment effect of $A$ on $Y$, one might use two of the groups enumerated in Figure 1-b (Standard RCT), yet if they only collect data concerning the groups in Figure 1-a (factorial experiment), this inference cannot be carried out unless one is willing to extrapolate from the factorial experiment to the standard one.
@@
a) 2-Treatment groups(Factorial experiment)$A=1;B=1$$A=1;B=0$$A=0;B=1$$A=0;B=0$b) 1-Treatment groups(Standard RCT)$A=1$$A=0$$B=1$$B=0$c) No-treatment group (Observational data)No intervention
@@
Figure 1: Study of individuals affected by covid-19 in a hospital. The outcome $Y$ represents their survival after two months. The two binary treatments $A$ and $B$ represent respectively the drug hydroxychloroquine and admission to the Intensive Care Unit (ICU). All the possibilities of treatment and control (interventions) are enumerated. In column (a), there are four possibilities of interventions at the same time on both $A$ and $B$. The set of those four cases is called here 2x2 factorial experiment. In column (b), there are four cases of single-treatment interventions on $A$ or $B$, which constitute a standard randomized controlled trial. Finally, in column (c), there is only one case when there is no active intervention. This is the case of observational data. It is important to emphasize that the observational data case, where $A$ and $B$ can be confounded with $Y$, cannot be mistaken for either the case $A=0;B=0$ in the factorial experiment because in this case the variables are not confounded, or the cases $A=0$ and $B=0$ in the standard RCT, because for $A=0$, $B$ might be confounded, and vice-versa.
@@
To make things more concrete, consider one has incomplete confounded data. For example, suppose one only has observational data and there is no experimental data. In such a world, without interventions, both hydroxychloroquine and ICU admission are confounded with $Y$ by unobserved confounders $U_{AY}$ and $U_{BY}$. This situation is represented by DAG (c) in Figure 2. Suppose one is interested in estimating the average treatment effect of hydroxychloroquine on COVID-19 survival (Figure 2-b). Because the only data they have is observational, and possible confounders are not observed, it is impossible to extrapolate from (c) to (b) without major assumptions. In other words, the only way we could estimate the effect of $A$ on $Y$, in this case, would be by implementing a standard RCT or trying to observe the confounders. Otherwise, any attempt to answer a causal question about Figure 2-b using data on Figure 2-c would be unproductive. This is the classical omitted variable bias, the known fact that an attempt to estimate causal effects from observational data with unobserved confounders will be biased.
@@
The converse problem of trying to answer an observational question using a standard RCT, while not widely known, has no solution either in this case. For example, suppose one runs an R.C.T. and assigns people to two randomized groups where they received hydroxychloroquine ($A=1$) or not ($A=0$). In this case, one can immediately answer questions regarding the average treatment effect of $A$ on $Y$. But if one intends to know observational facts about the incidence of the disease in the population, for instance, what percentage of COVID-19 patients survived in one hospital, the experimental data would not be sufficient. In other words, one cannot extrapolate from the single-treatment experiment represented in Figure 2-b, where there is intervention on $A$, and $B$ remains confounded with $Y$, to the observational world represented in Figure 2-c, where both $A$ and $B$ are confounded with $Y$.
@@
Finally, suppose that rather than implementing a randomized controlled trial with interventions only on $A$, one carries out a factorial experiment, and now they have four groups with interventions on $A$ and $B$ at the same time. This situation is represented by Figure 2-a. The problem now has changed to answering questions about single-treatment effects of $A$ (Figure 2-b), using the factorial data (Figure 2-a). And the answer, analogously to the previous case – as it will be shown – is negative. This result might be puzzling, given that we indeed had interventions on $A$ in the factorial experiment, but to answer the causal question about the single-treatment effect of $A$, one is implicitly assuming a world where $B$ remains confounded with $Y$. This result becomes still more puzzling when one realizes that even if they possessed at the same time factorial-experimental and observational data at the same time, the single-treatment effect of $A$ would not be immediately point-identifiable without other assumptions, as I will show. This result suggests one should be aware of potential problems of answering causal questions through cross-counterfactual world data.
@@
Figure 2: Graphs representing experimental/observational settings of a study trying to assess the effect of hydroxychloroquine ($A$) on COVID-19 survival ($Y$). $B$ denotes admission or not to the ICU. In the graphs, direct arrows denote causality direction. $U_{AY}$ and $U_{BY}$ are potential unobserved variables, and for that reason, we employed dashed arrows. Variables in boxes denote interventions, for example, fixing the value $a$ to the variable $A$ by an intervention. When an intervention happens, any arrow points to those intervened variables are immediately truncated. I consider three cases here. In case (a), representing a factorial experiment, there are two interventions on $B$ and $Y$, so confounding between $A$ and $Y$ and $B$ and $Y$ are removed. In case (b), a standard RCT, there is only one intervention setting $A$ to a. Finally, case (c) represents an observational setting, and both $A$ and $B$ are confounded with $Y$. Without major assumptions, extrapolations using data from one case to answer questions about other cases are not justified.
@@
In this paper, I assess when extrapolation from one case to another with different quantities of intervened variables is justified, and I develop a principled approach to estimate best- and worst-case scenarios for any estimand of interest in nonparametric settings. The approach consists of the following steps. First, researchers must specify the different worlds they are studying: observational, 1-treatment, …, and n-treatments as well as the causal structure they are willing to justify. Secondly, they have to consider potential reasonable functional assumptions for the scenario, for instance, additive separability, monotonicity, or linearity. Finally, they are able to derive point- or partial-identification results, using the algorithm derived in Duarte et al., (2023). Researchers introduce their data, set their causal-structural assumptions, ask a causal query, and our solution immediately returns a range of possibilities in terms of average effects. This range is guaranteed to be sharp in the sense that it cannot be improved given the extant information. Finally, we showed which assumptions – functional or structural – make the extrapolation possible in terms of point identification. If researchers are willing to justify some functional forms – for instance, linearity – then a point-identified solution might exist.
@@
This paper has the following structure. In section 2, I illustrate the problem by elaborating on an experiment with two treatments. I show how an estimated causal effect is biased if some of those assumptions fail. I demonstrate the types of assumptions required for the identification of single-treatment effects using factorial experiments, then show how current approaches are insufficient for identification when these are violated. In section 3, I derive new partial-identification results to address these shortcomings. Finally, in section 4, I apply these techniques to two simulated example sand demonstrate how they can be extended to sensitivity analyses that probe how results vary depending on key unmeasured quantities.
2 Identification problem
@@ 2.1 Non-identifiability result
Here I discuss how to derive identifiability when there is an arbitrary set of observational and experimental data and a specific causal query. In other words, one researcher might have available information on a factorial experiment, denoted here by the probability of an outcome given two treatments at the same time, $\Pr(Y_{a,b})$. For the purpose of notation, we use uppercase letters ($A,B,Y$) to represent random variables and lowercase letters ($a,b,y$) to represent their instantiations. Sometimes, this researcher might have observational data as well, denoted here by $\Pr(Y,A,B)$. In the paper, I will be assuming the standard regular assumptions of SUTVA and positivity (Imbens and Rubin,, 2015). The problem here is whether it is possible to identify single-treatment quantities using these data. In other words, one desires to know whether it is possible to identify $\Pr(Y_{a})$, using information on $\Pr(Y_{a,b})$ and $\Pr(Y,A,B)$.
Figure 3: A nonparametric structural graph illustrating a scenario where both $A$ and $B$ cause $Y$. All three variables, $A$, $B$, and $Y$, are influenced by unobserved confounders simultaneously. With the exception of cases where structural assumptions are explicitly relaxed, all the results presented in the paper assume this graph.
$B$$Y$$A$
In the context of the nonparametric scenario depicted in Figure 3, the answer is negative. Factorial experiments, whether conducted alone or alongside observational data, are insufficient for addressing questions regarding single-treatment effects. Demonstrating non-identifiability necessitates the construction of counterexamples wherein at least two distinct models yield identical data but possess estimands with differing values. In the appendix, I provide such counterexamples, illustrating this outcome both in a general setting and under the assumption of monotonic effects for both treatments. Intuitively, this arises because inquiries into a causal domain involving only one intervention cannot be directly resolved using information from causal domains featuring multiple interventions or none at all. Formally, this conclusion is derived from the following theorem:
@@
Consider treatments $A$ and $B$, along with an outcome $Y$. Assume that both $A$ and $B$ independently cause $Y$, without causing each other. Additionally, suppose the effects of both $A$ and $B$ on $Y$ are monotonic, meaning $\Pr(Y_{a_{0}}=y_{1},Y_{a_{1}}=y_{0})=\Pr(Y_{b_{0}}=y_{1},Y_{b_{1}}=y_{0})=0$. Under these conditions, without imposing any other assumptions, such as absence of confounding between $A$, $B$, and $Y$, the probability distribution $\Pr(Y_{a})$, representing the outcome under different levels of treatment $A$, is not identifiable solely from the joint distribution $\Pr(Y_{a,b})$ and $\Pr(Y,A,B)$. Furthermore, the Average Treatment Effect (ATE), defined as the difference in probabilities between different treatment levels, $\Pr(Y_{a}=y)-\Pr(Y_{a^{\prime}}=y)$, remains unidentifiable. Finally, these quantities are not identifiable even if we relax the assumption of monotonicity or if we only have observational or factorial data alone.
@@
This finding is sufficient to establish the non-identifiability of the estimand given the problem. Consequently, any endeavor to estimate the single-treatment quantity using data from two treatments and observational data without substantial assumptions will yield biased results. I term this bias F-Bias.
@@
The presence of $F$-bias raises the question of whether other functional assumptions, besides monotonicity, would lead to point-identification in those settings. The answer is affirmative, as the next subsection will demonstrate regarding additive separability (no interactions). However, these scenarios require both factorial and observational data. While the Average Treatment Effect (ATE) is point-identifiable, the expected value of isolated single-treatment estimands, such as $\operatorname{\mathbb{E}}(Y_{a})$, is not. Formal results are included in Theorem B.2. Furthermore, the identification of the ATE is no longer valid if either observational or factorial data are absent (Theorem B.3). Therefore, these scenarios will almost surely cause $F$-bias.

### v3

# Extrapolating Single-Treatment Effects Out of Factorial Experiments

@@ Abstract
Despite their cost, randomized controlled trials (RCTs) are widely regarded as gold-standard evidence in disciplines ranging from social science to medicine. In recent decades, researchers have increasingly sought to reduce the resource burden of repeated RCTs with factorial designs that simultaneously test multiple hypotheses, e.g. experiments that evaluate the effects of many medications or products simultaneously. This paper shows that when multiple interventions are randomized in experiments, the extrapolation of single-treatment effects outside the experimental setting is generally not identified without strong assumptions, even under otherwise realistic conditions. This happens because single-treatment effects involve a counterfactual world with a single focal intervention, allowing other variables to take their natural values (which may be confounded or modified by the focal intervention). In contrast, observational studies and factorial experiments provide information about potential-outcome distributions with zero and multiple interventions, respectively. Therefore, unobserved confounding affects not only the extrapolation from observational to experimental settings but also the extrapolation from factorial experiments to single-treatment settings. In this paper, I formalize sufficient conditions for the identifiability of those isolated quantities. Researchers who rely on this type of design need to justify either functional and/or structural assumptions. Finally, I develop nonparametric sharp bounds—i.e., maximally informative best-/worst-case estimates consistent with limited RCT data—that show when extrapolations about effect signs are empirically justified. These new results are illustrated with simulated data.
Keywords: Factorial Experiment; Partial identification; Transportability; External Validity; Causal inference;
@@ 1 Introduction
Randomized controlled trials are now widely considered the gold standard in empirical research. Researchers have used them to assess effects ranging from medication to economic development and discrimination. But their logistical complexity and cost are compounded by scientists’ common need to test many hypotheses. As a result, scholars turn to factorial experiments, which evaluate multiple hypotheses simultaneously. Papers using this general design number in the tens of thousands [^1] Recent theoretical results were reported by Dasgupta et al. (2015); Lee et al. (2020); Lee and Shpitser (2020); Egami and Imai (2018); De la Cuesta et al. (2022); Zhao and Ding (2022); Blackwell and Pashley (2023); Pashley and Bind (2023). and variants like conjoint analysis are an increasingly dominant approach in market research and survey experiments. Among the advantages of factorial designs, statisticians believe they provide greater efficiency because they allow the evaluation of factors with only one-quarter of the data that otherwise would be necessary, and greater comprehensiveness, as not only a few factors are analyzed but also their interactions (Fisher, 1949, p. 113). Recently, Nobel Laureates in Economics Esther Duflo and Michael Kremer argue, “because they significantly reduce costs, cross-cutting different treatments [i.e., factorial design] has proved very important in allowing for the recent wave of randomized evaluations in development economics” (Duflo et al., 2007). Randomization is often understood to allow scientists to infer “model-free,” or nonparametric, estimates of average treatment effects.
@@
However, factorial experiments do not generally support extrapolation to single-treatment effects without additional assumptions. For instance, while the average factorial effect for the target population is identifiable, a corresponding average single-treatment effect is only identifiable under additional assumptions. Intuitively, this is because factorial experiments present a setting where multiple treatments are implemented at the same time. However, when researchers are interested in single-treatment quantities, they are devising environments where one treatment is manipulated but all the others might be confounded with the outcome, i.e. other variables might be causing the treatment variable and the outcome at the same time. As those cases can only be assessed in trials with single treatments, any extrapolation requires additional identifying assumptions; in the fully nonparametric case, identification generally fails.
@@
Consider the following example. Suppose one is investigating individuals admitted to a hospital with symptoms of COVID-19. Let a binary $Y$ denote survival after 2 months. Assume $A$ denotes drug A, a drug often employed to treat lupus, and $B$ denotes admission or not to the Intensive Care Unit. Both are binary variables. One would like to know not only the joint effect of the two treatments, $A$ and $B$, on the outcome $Y$ but also their effects in isolation, for instance, the single-treatment effect of $A$ on $Y$ or $B$ on $Y$. If one designs a study, there are 9 cases we could consider: 1) 4 groups with 2-treatment interventions, A=B=1, A=0/B=1, A=1/B=0, and A=B=0; 2) 4 groups with 1-treatment interventions, A=0, A=1, B=0, B=1; 3) 1 group with no intervention, i.e. observational data. [^2] The observational group is different from the treatment group with A=0;B=0, because in the observational group, individuals might eventually receive $A$ and $B$ treatments given realizations of unobserved confounders. They are only the same in the particular situation where naturally no individual receives any treatment, which is the case of some program evaluations. All those cases are represented in Figure 1. The first case is exactly the case of a factorial experiment, while the second case is a single-armed randomized controlled trial (RCT), and the third represents observational data. Notice that observational data is not the same thing as the case $A=0;B=0$ in the factorial experiment, because $A$ and $B$ might be confounded with $Y$ in the no-treatment group. Indeed those cases are not even similar to groups $B=0$ and $A=0$ in the factorial RCT, because if one intervenes on $A$, $B$ might still be confounded, and vice-versa, if one intervenes on $B$. Therefore, if one intends to draw inferences on the average treatment effect of $A$ on $Y$, one might use two of the groups enumerated in Figure 1-b (single-armed RCT), yet if they only collect data concerning the groups in Figure 1-a (factorial experiment), this inference is not directly supported unless one is willing to extrapolate from the factorial experiment to the single-armed one.
@@
a) 2-Treatment groups(Factorial experiment)A=1;B=1A=1;B=1A=1;B=0A=1;B=0A=0;B=1A=0;B=1A=0;B=0A=0;B=0b) 1-Treatment groups(Single-treatment RCT)A=1A=1A=0A=0B=1B=1B=0B=0c) No-treatment group (Observational data)No intervention
@@
Figure 1: Study of individuals affected by COVID-19 in a hospital. The two binary treatments $A$ and $B$ represent respectively the drug A and admission to the Intensive Care Unit (ICU). All the possibilities of treatment and control (interventions) are enumerated. In column (a), a 2x2 factorial experiment, there are four possibilities of interventions at the same time on both $A$ and $B$. In column (b), there are four cases of single-treatment interventions on $A$ or $B$, which constitute a single-armed RCT. Finally, in column (c), there is only one case when there is no active intervention. This is the case of observational data.
@@
To make things more concrete, consider a setting with incomplete confounded data. For example, suppose one only has observational data and there is no experimental data. In such a world, without interventions, both drug A and ICU admission are confounded with $Y$ by unobserved confounders $U_{AY}$ and $U_{BY}$. This situation is represented by DAG (c) in Figure 2. Suppose one is interested in estimating the average treatment effect of drug A on COVID-19 survival. The only available data is observational, and possible confounders are not observed, it is not identifiable to extrapolate from (c) to (b) without strong assumptions. In other words, the only way we could estimate the effect of $A$ on $Y$, in this case, would be by implementing a single-treatment RCT or trying to observe the confounders. Otherwise, any attempt to answer a causal question about Figure 2-b using data on Figure 2-c would not be identifiable. This is the classical omitted variable bias, the known fact that an attempt to estimate causal effects from observational data with unobserved confounders will be biased.
@@
The converse problem of trying to answer an observational question using a single-treatment RCT, while not widely considered, is not identifiable either in this case. For example, suppose one runs an RCT and assigns people to two groups where they received drug A ($A=1$) or not ($A=0$). In this case, one can answer questions regarding the average treatment effect of $A$ on $Y$. But if one intends to know observational facts about the incidence of the disease in the population, for instance, what percentage of COVID-19 patients survived in one hospital, the experimental data would not be sufficient, or the survival rate for those who were admitted to the ICU by instruction of the physician. In other words, it is not identifiable to extrapolate from the single-treatment experiment represented in Figure 2-b, where there is intervention on $A$, and $B$ remains confounded with $Y$, to the observational world represented in Figure 2-c, where both $A$ and $B$ are confounded with $Y$.
@@
Finally, suppose that rather than implementing an RCT with interventions only on $A$, one carries out a factorial experiment and now has four groups with interventions on $A$ and $B$ at the same time. That is a common scenario for researchers who are not exactly interested in the effect of those treatments at the same time but implement them anyway for for cost-saving reasons. The problem now has changed to answering questions about single-treatment effects of $A$ (Figure 2-b), using the factorial data (Figure 2-a). And the answer, analogously to the previous case – as shown below – is generally negative. This result might be puzzling, given that we indeed had interventions on $A$ in the factorial experiment, but to answer the causal question about the single-treatment effect of $A$, one is implicitly assuming a world where $B$ remains confounded with $Y$. This result becomes still more puzzling when one realizes that even if they possessed at the same time factorial-experimental and observational data, the single-treatment effect of $A$ would not be point-identifiable without other assumptions, as shown below. This result suggests one should be aware of potential problems of answering causal questions through cross-counterfactual world data.
@@
Figure 2: Graphs representing experimental/observational settings of a study trying to assess the effect of drug A ($A$) on COVID-19 survival ($Y$). $B$ denotes admission or not to the ICU. In the graphs, direct arrows denote causality direction. $U_{AY}$ and $U_{BY}$ are potential unobserved variables, and for that reason, we employed dashed arrows. Any direct effect of $A$ on $B$ or vice-versa is not considered here (no mediation). Variables in boxes denote interventions, for example, fixing the value $a$ to the variable AA by an intervention. When an intervention happens, any arrow points to those intervened variables are immediately truncated. I consider three cases here. In case (a), representing a factorial experiment, there are two interventions on $B$ and $Y$, so confounding between $A$ and $Y$ and $B$ and $Y$ are removed. In case (b), a single-treatment RCT, there is only one intervention setting AA to a. Finally, case (c) represents an observational setting, and both $A$ and $B$ are confounded with $Y$. Without major assumptions, extrapolations using data from one case to answer questions about other cases are not supported.
@@
In this paper, I assess when extrapolation from one case to another with different quantities of intervened variables is justified, and I develop a principled approach to estimate best- and worst-case scenarios for any estimand of interest in nonparametric settings. The approach consists of the following steps. First, researchers should specify the different worlds they are studying: observational, 1-treatment, …, and n-treatments as well as the causal structure they are willing to justify. Secondly, they consider potential reasonable functional assumptions for the scenario, for instance, additive separability, monotonicity, or linearity. Finally, they are able to derive point- or partial-identification results.
The problem of generalizability from experimental settings to target populations has a long history in statistics and causal inference. Fisher (1949) recognized those concerns and emphasized that treatment comparisons obtained under particular experimental conditions need not extend to other settings, and argued that treatment–environment interactions are the appropriate objects for assessing the stability of experimental conclusions. Within Fisher’s framework, however, the generalizability problem was understood as being resolved once these interactions were estimated: given sufficient factorial variation across environments, treatment effects were taken to be fully characterized by their interaction structure. When reformulated in modern causal terms, this view is incomplete. Even with complete knowledge of treatment–environment interactions on the experimental support, there is no implied extrapolation from those environments to a target population, and hence no uniquely defined single-treatment effect for real-world deployment. [^3] Knowing all interactions tells us how treatments behave where we experimented, but not how to aggregate or transport those effects to the environment in which a policy is actually applied. By making the target estimand explicit, modern causal formalizations reveal that the absence of an extrapolation rule constitutes an identification problem rather than an estimation problem. The contribution of this paper is to address this gap by formalizing generalizability in causal terms and by providing partial identification bounds for target-population treatment effects under transparent and weak assumptions.
@@
This paper has the following structure. In section 2, I illustrate the problem by elaborating on an experiment with two treatments. I show how an estimated causal effect can be biased if some of those assumptions fail. I demonstrate the types of assumptions required for the identification of single-treatment effects using factorial experiments, then show how current approaches are insufficient for identification when these are violated. In section 3, I derive new partial-identification results to address these shortcomings. Finally, in section 4, I apply these techniques to two simulated examples and demonstrate how they can be extended to sensitivity analyses that probe how results vary depending on key unmeasured quantities.
2 Extrapolation problem
@@ 2.1 Non-identifiability result
Here I study whether single-treatment effects can be identified when there is an arbitrary set of observational and experimental data and a specific causal query. In other words, one researcher might have available information on a factorial experiment, and maybe observational data as well, and they want to know whether the single-treatment effect can be extrapolated from those data. The answer is generally negative, if no strong assumptions are made, as I show below. On the other hand, if one is willing to make strong structural or functional assumptions, then identification might be possible.
For the purpose of notation, we use uppercase letters ($A,B,Y$) to represent random variables and lowercase letters ($a,b,y$) to represent their instantiations. Sometimes, this researcher might have observational data as well, denoted here by $\Pr(Y,A,B)$. In the paper, I will be assuming the standard regular assumptions of SUTVA and positivity (Imbens and Rubin, 2015). The notation of potential outcomes is used throughout the paper. $Y_{a,b}$ denotes the potential outcome $Y$ when treatments $A$ and $B$ are set to $a$ and $b$, respectively. Similarly, $Y_{a}$ denotes the potential outcome $Y$ when treatment $A$ is set to $a$, and $B$ takes its natural value. The problem here is whether it is possible to identify single-treatment quantities using these data. In other words, one desires to know whether it is possible to identify $\Pr(Y_{a})$, using information on $\Pr(Y_{a,b})$ and $\Pr(Y,A,B)$.
Figure 3: A nonparametric structural graph illustrating a scenario where both $A$ and $B$ cause $Y$. All three variables, AA, $B$, and $Y$, are influenced by unobserved confounders simultaneously. With the exception of cases where structural assumptions are explicitly relaxed, all the results presented in the paper assume this graph.
BBYYAA
In the context of the nonparametric scenario depicted in Figure 3, identification does not generally hold. Factorial experiments, whether conducted alone or alongside observational data, are insufficient for addressing questions regarding single-treatment effects. Demonstrating non-identifiability necessitates the construction of counterexamples wherein at least two distinct models yield identical data but possess estimands with differing values. In the appendix, I provide such counterexamples, illustrating this outcome both in a general setting and under the assumption of monotonic effects for both treatments. Intuitively, this arises because inquiries into a causal domain involving only one intervention cannot be directly resolved using information from causal domains featuring multiple interventions or none at all. Formally, this conclusion is derived from the following theorem:
@@
Consider treatments $A$ and $B$, along with an outcome $Y$. Assume that both $A$ and $B$ independently cause $Y$, without causing each other. Additionally, suppose the effects of both $A$ and $B$ on $Y$ are monotonic, meaning $\Pr(Y_{a_{0}}=y_{1},Y_{a_{1}}=y_{0})=\Pr(Y_{b_{0}}=y_{1},Y_{b_{1}}=y_{0})=0$. Under these conditions, without imposing any other assumptions, such as absence of confounding between $A$, $B$, and $Y$, the probability distribution $\Pr(Y_{a})$, representing the outcome under different levels of treatment $A$, is not identifiable solely from the joint distribution $\Pr(Y_{a,b})$ and $\Pr(Y,A,B)$. Furthermore, the ATE, defined as the difference in probabilities between different treatment levels, $\Pr(Y_{a}=y)-\Pr(Y_{a^{\prime}}=y)$, remains unidentifiable. Finally, these quantities are not identifiable even if we relax the assumption of monotonicity or if we only have observational or factorial data alone.
@@
This finding is sufficient to establish the non-identifiability of the estimand given the problem. Consequently, attempts to estimate the single-treatment quantity using data from two treatments and observational data without substantial assumptions can yield biased results. I term this bias F-Bias.
@@
The presence of $F$-bias raises the question of whether other functional assumptions, besides monotonicity, would lead to point-identification in those settings. The answer is affirmative, as the next subsection will demonstrate regarding additive separability (no interactions). However, these scenarios require both factorial and observational data. While the ATE is point-identifiable, the expected value of isolated single-treatment estimands, such as $\operatorname{\mathbb{E}}(Y_{a})$, is not. Formal results are included in Theorem B.2. Furthermore, the identification of the ATE is no longer valid if either observational or factorial data are absent (Theorem B.3). Therefore, these scenarios will typically cause $F$-bias.
It should be emphasized that the problem here is one of transportability (of extrapolation). For the population defined in the experimental design—which we denote by $Y^{F}_{a}$ to distinguish it from the population of interest— the quantity $\Pr(Y^{F}_{a})$ is identifiable from the experimental data $\Pr(Y^{F}_{a,b})$ by the law of total probability: $\Pr(Y^{F}_{a})=\sum_{b}\Pr(Y^{F}_{a,b})\Pr(B^{F}=b)$. However, the problem here is to identify $\Pr(Y_{a})$ for the target population, which might be different from the experimental one. In other words, the problem is to extrapolate from $\Pr(Y^{F}_{a,b})$ and $\Pr(Y,A,B)$ to $\Pr(Y_{a})$. A common practice in this scenario is to state that $\Pr(Y^{F}_{a})$ is just a second-best estimand, to recognize that it lacks external validity, and to hope that it is close enough to $\Pr(Y_{a})$. However, as I show, even under otherwise realistic conditions, this extrapolation is not generally supported without strong assumptions. In this paper, we provide precise conditions under which this extrapolation is justified.